You bought an apartment which wan in a 20-year mortgage of 10,000,000 from the bank ABC. The bank ABC asked you pay back the loan by monthly payment on 3.6% nomonal interest rate with 3-year grace(pay interest only). You have stayed with the bank ABC for 5 yesrd and your Minimum Attractive Rate of Return(MARR) is 0.75% per month

(1)What is the monthly payment that you are asked to make now

(2)what is the interest you have paid so far

(3)what is the remaining principal
Answer #1
A Loan Amount $10,000,000
B Monthly interest =(3.6/12)% 0.30%
C=A*B Monthly Payment for first three years $30,000
1 Amount you are asked to pay now (per month) $30,000
D=C*3*12 Total payment in three years $1,080,000
Pv Principal Balance at the end of 3 years $10,000,000
Nper Balance payment period in months=17*12 204
Rate Monthly interest rate 0.30%
PMT Monthly payment from year 4 onwards $65,611.39 (Using PMT function of excel with Rate =0.3%, Nper=204,Pv=-10000000)
E=PMT*24 Payment made in balance 2 years(5-3) $1,574,673
F=D+E Total payment made in 5 years $2,654,673
Principal Balance at end of 5 years =Present Value of remaining payment for 15 years
Rate Interest rate 0.30%
Nper Number of months payment remaining after year5 180 (15*12)
Pmt Monthly payment $65,611.39
PV Principal Balance at end of 5 years = $9,115,181 (Using PV function of excel with Rate =0.3%, Nper=180,Pmt=-65611.39)
G Total Principal Paid $884,819 (10000000-9115181)
2 Interest paid so far =2654673-884819 $1,769,855
3 Amount of remaining Principal $9,115,181